OpenStudy (anonymous):

The equation: y=144x+621, where y represents the cost in dollars and x=0 represents the year 1980. Use the equation to find the cost of tuition and fees in 2002 and the year this cost will exceed $5250

OpenStudy (anonymous):

If x=0 represents 1980 and we assume that x=years, then 2002 will be x=22 isn't it? So y for x=22 will be y=144*22+621=3789 So, in 2002 the cost of fees is $3789 Now, in part "b" you want to know the yeat this cost will EXCEED 5250, and that's an inequation \[144\times x+621>y\] For y=5250 \[144\times x+621>5250\] \[x>\dfrac {5250-621} {144}\] That's aprox x>32.15 So, in the year "33" the fee costs will have exceeded the quantity, and, if 1980 goes for x=0, x=33 means 1980+33=2013 Hope you understood D:

OpenStudy (anonymous):

I understood part A but not so much on part B a bit confuse on the inequation

OpenStudy (anonymous):

ohhh never mind!! now that I look at it, the word EXCEED means bigger than 5250, so just replace y with 5250 and the orginal equation then find x

OpenStudy (anonymous):

cuz u put 144 times X, made me confuse a little lol :)

OpenStudy (anonymous):

Cool ;)